please help with this word problem. I need the equation to figure this out. thanks in advance. this website is great!
Ellen wishes to mix candy worth $1.90 per pound with candy worth $3.50 per pound to form 32 pounds of a mixture worth $2.80 per pound. How many pounds of the more expensive candy should she use?
thanks again,

Answer by Paul(988)

1.9x+3.5(32-x)=2.80(32)
1.9x+112-3.5x=89.6
-1.6x=-22.4
x=14
32-14=18
Hence the solution for $1.9 is 14 and for $3.5 is 18